# Learning About Structural Errors in Models of Complex Dynamical Systems

### v1

@@ 2.2.2 Enforcing Constraints
Sparsity is important to impose on error models to avoid adding parameters to an error model that have no or little measurable impact on the output $\hat{y}$ of the host model; that is, the first principle in modeling structural errors should be to do no harm that would result from unnecessary error model complexity. To impose sparsity on the solution of $\theta_{\textrm{I}}$, we aim to solve the optimization problem Schneider et al. (2022)
@@
where $\mathfrak{V}=\{\theta:|\theta_{\textrm{I}}|_{\ell_{1}}\leq\gamma\}$. The regularization parameters $\gamma$ and $\lambda$ can be determined via cross-validation. In practice, adding the $\ell_{0}$ constraint is achieved by thresholding the results from $\ell_{1}$-constrained optimization. The detailed algorithm was proposed in Schneider et al. (2022) and is summarized in A.
@@ 2.2.3 Accounting for Initial Conditions
Evaluation of $\mathcal{G}$ typically requires an initial condition, $X(0)$, when solving Eq. (5). When we observe ergodic systems over sufficiently long time scales, we can treat $\mathcal{G}$ as a stochastic map over a distribution of initial conditions Cleary et al. (2021). However, when we observe systems over shorter time scales (or if they are non-ergodic), it is often necessary to identify initial conditions that generated the observations. We highlight two basic approaches: (i) append $X(0)$ to the vector of learned parameters for $\mathcal{G}$, and (ii) implicitly identify $X(0)$ by applying data assimilation techniques. To achieve (i), we can minimize:
@@
where $\mathcal{DA}$ represents a data assimilation algorithm that performs state estimation given observations $y$ and assumed parameters $\theta$. While Eqs. (13) and (14) appear quite similar, the utilization of efficient state-estimation algorithms in (14) can help cope with sensitivities to initial conditions that may cause (13) to fail. Recent work has explored different variants of (14), often using gradient-based optimization approaches and sequential state-estimation schemes Levine and Stuart (2022); Ribera et al. (2022); Brajard et al. (2021); Chen et al. (2022).
@@ 3.1 Dictionary Learning
where $\theta_{\textrm{I}}=\{\alpha,\beta\}$ and $\phi_{j}(X;\beta_{j})$ denote user-specified, parametric basis functions that can, for example, include differential operators Brunton et al. (2016); Rudy et al. (2017); Schaeffer et al. (2013); Schaeffer (2017). In practice, it is difficult to know all suitable basis functions a priori, and thus it is common to include redundant basis functions in the dictionary. Basis functions can then be pruned based on data by imposing sparsity constraints on the coefficients $\alpha_{j}$. Such sparsity constraints have proven to be beneficial in the construction of data-driven models of dynamical systems Brunton et al. (2016); Rudy et al. (2017); Schaeffer et al. (2013); Schaeffer (2017); Schneider et al. (2022). They are also commonly used in compressed sensing Donoho (2006), where dictionary learning has been widely used.
@@ 3.2 Gaussian Processes
Another option of constructing an error model is via Gaussian processes (GPs) Williams and Rasmussen (2006), [^5] Here we use only the mean of the GP, and the methodology is simply a form of data-adapted regression; we are not using the uncertainty quantification that comes with GPs.
@@
Because of the similar forms of Eqs. (17) and (15), the GP shares similar shortcomings as dictionary learning when the kernel $\mathcal{K}$ is misspecified, even in the presence of hyper-parameter learning. In practice, a more sophisticated kernel $\mathcal{K}=\sum\limits_{i}\mathcal{K}_{i}$ is often constructed from some basic kernels $\mathcal{K}_{i}$ Hamzi and Owhadi (2021); Darcy et al. (2021); Lee et al. (2022). If a redundant set of basic kernels is used, sparsity constraints can be imposed in a similar way as in dictionary learning to prune the kernel set. A further limitation of using GPs is the computational cost, which grows exponentially with the dimension of $\mathcal{X}$. This pathology can be ameliorated by representing the GP as a linear combination of random Fourier features Rahimi and Recht (2008), which allows us to recast a GP as a dictionary-based approach in which the bases $\phi_{j}$ are drawn randomly from a special distribution known to reproduce a kernel of interest.
@@ 3.3 Neural Networks
where $\mathcal{NN}$ denotes a neural network and $\theta_{\textrm{I}}$ the coefficients (biases and weights) of the neural network. While neural networks are expressive and scalable, it is more difficult to enforce stability of a dynamical system with a neural network error model Brenowitz et al. (2020). This is mainly because the nonlinearity introduced by a neural network is often more difficult to analyze compared with dictionary learning, for which we explicitly specify basis functions and thus can avoid using those that lead to instability; it is also more difficult to analyze than GP based learning because the latter is easier to interpret, as the kernels are hand-picked and then tuned to data. In Section 2.2.2, we discuss a general approach to enhancing stability by enforcing energy constraints in the context of learning from indirect data.
@@ 3.4 Stochastic Extension
Fitting stochastic models to time-series data has been explored in some previous works Wilkinson (2009); Roberts and Stramer (2001, 2002); Papaspiliopoulos et al. (2013); a common problem when applying these methods is the inconsistency between data and the incremental structure of the Gaussian noise driving the model as the time step is approaching zero Zhang et al. (2005); Pavliotis and Stuart (2007); Papavasiliou et al. (2009); Boninsegna et al. (2018). A common practice to address this issue is the multi-scale use of data, e.g., via subsampling Zhang (2006); Pokern et al. (2009); Papaspiliopoulos et al. (2012); Pavliotis et al. (2012); Batz et al. (2018); Abdulle et al. (2021). Some previous works also explored Kramers–Moyal averaging with finite sampling rate correction Honisch and Friedrich (2011); Lade (2009); Callaham et al. (2021). On the other hand, fitting a discretized version of stochastic processes to time-series data has been explored using autoregressive models Neumaier and Schneider (2001); Schneider and Neumaier (2001); Arnold et al. (2013); Lu (2020). For some dynamical systems, the unresolved state has conditional (with respect to the resolved state) Gaussian statistics Chen and Majda (2016, 2018), and then fitting the stochastic models can be achieved using analytically derived likelihoods.
In the absence of the whole trajectories of time series, some recent works started to explore fitting stochastic models to statistics of time-series data Krumscheid et al. (2013, 2015); Kalliadasis et al. (2015); Schneider et al. (2021). Using time-averaged data to estimate linear SDEs has been studied for decades to account for climate variability Hasselmann (1976); Frankignoul and Hasselmann (1977); Penland and Magorian (1993); Schneider and Griffies (1999), and extension to nonlinear SDEs was discussed in Hasselmann (1988).
@@ 3.5 Representing Non-local Effects
In the case of neural networks, we can build spatially non-local closures with convolutional neural networks (CNNs); the Fourier neural operator (FNO) Li et al. (2020) or deep operator network (DeepONet) Lu et al. (2021) provide an extension to an operator limit. Similarly, GPs and random feature methods (a dictionary-based formulation of GPs) can be designed with spatially non-local vectorized inputs from $X(t)$. Recent theoretical work has also allowed these basic methods to be taken to a continuous operator limit Nelsen and Stuart (2021); Chen et al. (2021).
As an emerging topic in the context of data-driven modeling, some recent works have explored non-local diffusion Du et al. (2012); D’Elia and Gunzburger (2016); Du (2019); Pang et al. (2020) and spatially non-local modeling Zhou et al. (2021); Zhuang et al. (2021); Charalampopoulos and Sapsis (2021). In this work, we capture the spatially non-local dependence on $X$ via a data-driven convolution kernel:
@@
Note that hyper-parameters can be made state-dependent, so that $\theta_{\textrm{non-loc}}(X(t,r);\kappa)$; in this case, the additional unknowns $\kappa$ can be learned, appending them to $\theta_{\textrm{I}}$. Similarly, learning a nonlinear integral kernel has been discussed in Kovachki et al. (2023) and shown to be a continuous generalization of the transformer architecture Vaswani et al. (2017).
The form of non-local closure in (20) draws inspiration from a series of works about non-local modeling Du (2019), in which $\delta_{\textrm{loc}}$ corresponds to a local Laplace operator. Some mathematical foundations of non-local operators and calculus were summarized in Du (2019), and the connection to fractional differential operators was illustrated in Bucur and Valdinoci (2016).
@@
Non-locality in time (memory) is also important. Generically, any form of variable elimination or coarse-graining results in memory effects which require, at each current point in time, integration of the entire time-history from the initial condition up to the current time Zwanzig (2001). Such memory effects are undesirable as they lead to computational algorithms that scale poorly with respect to length of the time interval. Markovian models that encapsulate memory can be constructed, for example, by introducing a recurrent neural network Levine and Stuart (2022), or by the use of delay embedding Sauer et al. (1991); such Markovian models are more computationally expedient. Temporally non-local modeling has received significant recent attention Ma et al. (2018); Wang et al. (2020); Charalampopoulos and Sapsis (2021); Lin and Lu (2021). If diffusion/advection mechanisms are present in the resolved system, memory effects of any state variable would manifest themselves in the state variables as a spatial non-locality. For this reason non-local models with a flexible enough kernel could potentially be used to capture memory effects, without significantly increasing the computational costs.
@@ 4 Lorenz 96 Systems as Illustrative Examples
For our simulation studies of structural model error, we consider two variants of the celebrated Lorenz 96 model Lorenz (1996), described in what follows.
@@ 4.1.1 Multiscale Lorenz 96 Model
The Lorenz 96 multi-scale system Lorenz (1996) describes the evolution of a simplified atmospheric flow, which is periodic along latitude circles (space). It does so through one set of slow variables, $x_{k}$ ($k=1,\dots,K$), coupled to a set of fast variables, $z_{j,k}$ ($j=1,\dots,J$), whose indices label space coordinates:
@@
of the fast variables affecting the slow variables. To generate data, we work with the parameter choices $K=36$, $J=10$, and $F=b=10$ Lorenz (1996); Schneider et al. (2017). The choices of $h$ and $c$ are summarized in Subsection 4.2 for different cases.
@@ 4.2.1 Lorenz 96 Multi-scale Model
    For $c=10$ and $h=1$ in the multi-scale Lorenz 96 system, a dictionary-learning-based model is trained as local deterministic error model $\delta(X_{k})$ using direct data ($\{x_{k},hc\overline{z}_{k}\}$), and a fully-connected neural network is trained using indirect data (first and second moments of the slow variable $x_{k}$). For the indirect data in this example, we assume partial observations of the first eight slow variables and include cross-terms of second moments (i.e., $\mathbb{E}(x_{i}x_{j})$ for different $i$ and $j$). The results are presented in Figs. 2 and 3.
@@
    For $c=3$ and $h=10/3$ in the multi-scale Lorenz 96 system, the scale separation between fast and slow variables becomes smaller and thus leads to a more challenging case. In this case, a fully-connected neural network is trained as the local deterministic error model $\delta(X_{k})$ using either direct data ($\{x_{k},hc\overline{z}_{k}\}$) or indirect data (first to fourth moments of the slow variable $x_{k}$ and the autocorrelation of $x_{k}$). For the indirect data in this and the next examples, we enable the full observation of all 36 slow variables and preclude the use of all cross-terms of second to fourth moments. The results are presented in Figs. 4 and 5.
@@
To avoid the overfitting in Fig. 5, we calibrate a non-local error model, as discussed in Section 3.5. Compared to the results for the local error model, it can be seen in Fig. 6 that the invariant measure of the calibrated system agrees better with the true system, which indicates that the closure model $\delta(\cdot)$ in (24) with non-local effects represents a better closure for unresolved scales if there is a less clear scale separation between resolved and unresolved scales.
@@
We also explored the learning of stochastic error model for this example. Figure 7 presents the results of the calibrated system with an additive stochastic error model. Compared to the results for deterministic error models in Figs. 5 and 6, we can see that the invariant measure of the calibrated system agrees better with the true system in Fig. 7b. We further test the stochastic error model by also learning a state-dependent diffusion coefficient. As shown in Fig. 8b, the calibrated system achieves better agreement with the invariant measure of the true system, which confirms that increased flexibility in the stochastic error model can help achieve improved predictive performance via training against indirect data. It should be noted that Figs. 6 to 8 do not display a single converged invariant measure. This arises because the calibrated parameters vary across the ensemble; as a result, we obtain a family of structural error models that all fit the data with similar accuracy. We surmise that this is caused by the fact that the indirect data contain only partial information about the invariant measure. Nonetheless, the fits are all far superior to those obtained with the local deterministic model.

### v2

@@ 2.2.2 Enforcing Constraints
Sparsity is important to impose on error models to avoid adding parameters to an error model that have no or little measurable impact on the output $\hat{y}$ of the host model; that is, the first principle in modeling structural errors should be to do no harm that would result from unnecessary error model complexity. To impose sparsity on the solution of $\theta_{\textrm{I}}$, we aim to solve the optimization problem [74]
@@
where $\mathfrak{V}=\{\theta:|\theta_{\textrm{I}}|_{\ell_{1}}\leq\gamma\}$. The regularization parameters $\gamma$ and $\lambda$ can be determined via cross-validation. In practice, adding the $\ell_{0}$ constraint is achieved by thresholding the results from $\ell_{1}$-constrained optimization. The detailed algorithm was proposed in [74] and is summarized in A.
@@ 2.2.3 Accounting for Initial Conditions
Evaluation of $\mathcal{G}$ typically requires an initial condition, $X(0)$, when solving Eq. (5). When we observe ergodic systems over sufficiently long time scales, we can treat $\mathcal{G}$ as a stochastic map over a distribution of initial conditions [75]. However, when we observe systems over shorter time scales (or if they are non-ergodic), it is often necessary to identify initial conditions that generated the observations. We highlight two basic approaches: (i) append $X(0)$ to the vector of learned parameters for $\mathcal{G}$, and (ii) implicitly identify $X(0)$ by applying data assimilation techniques. To achieve (i), we can minimize:
@@
where $\mathcal{DA}$ represents a data assimilation algorithm that performs state estimation given observations $y$ and assumed parameters $\theta$. While Eqs. (13) and (14) appear quite similar, the utilization of efficient state-estimation algorithms in (14) can help cope with sensitivities to initial conditions that may cause (13) to fail. Recent work has explored different variants of (14), often using gradient-based optimization approaches and sequential state-estimation schemes [51, 76, 77, 58].
@@ 3.1 Dictionary Learning
where $\theta_{\textrm{I}}=\{\alpha,\beta\}$ and $\phi_{j}(X;\beta_{j})$ denote user-specified, parametric basis functions that can, for example, include differential operators [50, 29, 28, 30]. In practice, it is difficult to know all suitable basis functions a priori, and thus it is common to include redundant basis functions in the dictionary. Basis functions can then be pruned based on data by imposing sparsity constraints on the coefficients $\alpha_{j}$. Such sparsity constraints have proven to be beneficial in the construction of data-driven models of dynamical systems [50, 29, 28, 30, 74]. They are also commonly used in compressed sensing [78], where dictionary learning has been widely used.
@@ 3.2 Gaussian Processes
Another option of constructing an error model is via Gaussian processes (GPs) [79], [^5] Here we use only the mean of the GP, and the methodology is simply a form of data-adapted regression; we are not using the uncertainty quantification that comes with GPs.
@@
Because of the similar forms of Eqs. (17) and (15), the GP shares similar shortcomings as dictionary learning when the kernel $\mathcal{K}$ is misspecified, even in the presence of hyper-parameter learning. In practice, a more sophisticated kernel $\mathcal{K}=\sum\limits_{i}\mathcal{K}_{i}$ is often constructed from some basic kernels $\mathcal{K}_{i}$ [80, 81, 82]. If a redundant set of basic kernels is used, sparsity constraints can be imposed in a similar way as in dictionary learning to prune the kernel set. A further limitation of using GPs is the computational cost, which grows exponentially with the dimension of $\mathcal{X}$. This pathology can be ameliorated by representing the GP as a linear combination of random Fourier features [83], which allows us to recast a GP as a dictionary-based approach in which the bases $\phi_{j}$ are drawn randomly from a special distribution known to reproduce a kernel of interest.
@@ 3.3 Neural Networks
where $\mathcal{NN}$ denotes a neural network and $\theta_{\textrm{I}}$ the coefficients (biases and weights) of the neural network. While neural networks are expressive and scalable, it is more difficult to enforce stability of a dynamical system with a neural network error model [60]. This is mainly because the nonlinearity introduced by a neural network is often more difficult to analyze compared with dictionary learning, for which we explicitly specify basis functions and thus can avoid using those that lead to instability; it is also more difficult to analyze than GP based learning because the latter is easier to interpret, as the kernels are hand-picked and then tuned to data. In Section 2.2.2, we discuss a general approach to enhancing stability by enforcing energy constraints in the context of learning from indirect data.
@@ 3.4 Stochastic Extension
Fitting stochastic models to time-series data has been explored in some previous works [84, 85, 86, 87]; a common problem when applying these methods is the inconsistency between data and the incremental structure of the Gaussian noise driving the model as the time step is approaching zero [88, 89, 90, 91]. A common practice to address this issue is the multi-scale use of data, e.g., via subsampling [92, 66, 93, 94, 95, 96]. Some previous works also explored Kramers–Moyal averaging with finite sampling rate correction [97, 98, 99]. On the other hand, fitting a discretized version of stochastic processes to time-series data has been explored using autoregressive models [100, 101, 102, 103]. For some dynamical systems, the unresolved state has conditional (with respect to the resolved state) Gaussian statistics [104, 105], and then fitting the stochastic models can be achieved using analytically derived likelihoods.
In the absence of the whole trajectories of time series, some recent works started to explore fitting stochastic models to statistics of time-series data [106, 107, 108, 52]. Using time-averaged data to estimate linear SDEs has been studied for decades to account for climate variability [109, 110, 111, 112], and extension to nonlinear SDEs was discussed in [113].
@@ 3.5 Representing Non-local Effects
In the case of neural networks, we can build spatially non-local closures with convolutional neural networks (CNNs); the Fourier neural operator (FNO) [114] or deep operator network (DeepONet) [115] provide an extension to an operator limit. Similarly, GPs and random feature methods (a dictionary-based formulation of GPs) can be designed with spatially non-local vectorized inputs from $X(t)$. Recent theoretical work has also allowed these basic methods to be taken to a continuous operator limit [116, 117].
As an emerging topic in the context of data-driven modeling, some recent works have explored non-local diffusion [43, 44, 118, 45] and spatially non-local modeling [41, 42, 49]. In this work, we capture the spatially non-local dependence on $X$ via a data-driven convolution kernel:
@@
Note that hyper-parameters can be made state-dependent, so that $\theta_{\textrm{non-loc}}(X(t,r);\kappa)$; in this case, the additional unknowns $\kappa$ can be learned, appending them to $\theta_{\textrm{I}}$. Similarly, learning a nonlinear integral kernel has been discussed in [119] and shown to be a continuous generalization of the transformer architecture [120].
The form of non-local closure in (20) draws inspiration from a series of works about non-local modeling [118], in which $\delta_{\textrm{loc}}$ corresponds to a local Laplace operator. Some mathematical foundations of non-local operators and calculus were summarized in [118], and the connection to fractional differential operators was illustrated in [121].
@@
Non-locality in time (memory) is also important. Generically, any form of variable elimination or coarse-graining results in memory effects which require, at each current point in time, integration of the entire time-history from the initial condition up to the current time [37]. Such memory effects are undesirable as they lead to computational algorithms that scale poorly with respect to length of the time interval. Markovian models that encapsulate memory can be constructed, for example, by introducing a recurrent neural network [51], or by the use of delay embedding [122]; such Markovian models are more computationally expedient. Temporally non-local modeling has received significant recent attention [46, 47, 49, 48]. If diffusion/advection mechanisms are present in the resolved system, memory effects of any state variable would manifest themselves in the state variables as a spatial non-locality. For this reason non-local models with a flexible enough kernel could potentially be used to capture memory effects, without significantly increasing the computational costs.
@@ 4 Lorenz 96 Systems as Illustrative Examples
For our simulation studies of structural model error, we consider two variants of the celebrated Lorenz 96 model [54], described in what follows.
@@ 4.1.1 Multiscale Lorenz 96 Model
The Lorenz 96 multi-scale system [54] describes the evolution of a simplified atmospheric flow, which is periodic along latitude circles (space). It does so through one set of slow variables, $x_{k}$ ($k=1,\dots,K$), coupled to a set of fast variables, $z_{j,k}$ ($j=1,\dots,J$), whose indices label space coordinates:
@@
of the fast variables affecting the slow variables. To generate data, we work with the parameter choices $K=36$, $J=10$, and $F=b=10$ [54, 65]. The choices of $h$ and $c$ are summarized in Subsection 4.2 for different cases.
@@ 4.2 Numerical Results for Lorenz Models
Here, we investigate the long-term behavior of the trained models for the Lorenz systems via the invariant measure. For the numerical examples, the invariant measure corresponds to the probability density function of a system state variable, when simulating the dynamical system for a long enough time so that the probability density function does not vary. For the numerical examples with indirect data, the invariant measure is employed as a qualitative indicator for the long-term performance of the trained models. Considering that the indirect data only contain partial information about the true invariant measure, the quantitative comparison of the ensemble mean for the invariant measures may lead to over-interpretation of the results; thus, we only present the ensemble of invariant measures.
@@ 4.2.1 Lorenz 96 Multi-scale Model
    For $c=10$ and $h=1$ in the multi-scale Lorenz 96 system, a fully-connected neural network is trained as a local deterministic error model $\delta(X_{k})$ using direct data ($\{x_{k},hc\overline{z}_{k}\}$), and a dictionary-learning-based model is trained using indirect data (first and second moments of the slow variable $x_{k}$). The direct data comprise 36000 data points. For the indirect data in this example, we assume partial observations of the first eight slow variables and include cross-terms of second moments (i.e., $\mathbb{E}(x_{i}x_{j})$ for different $i$ and $j$). The indirect data comprise 44 data points. The results are presented in Figs. 2 and 3.
@@
    For $c=3$ and $h=10/3$ in the multi-scale Lorenz 96 system, the scale separation between fast and slow variables becomes smaller and thus leads to a more challenging case. In this case, a fully-connected neural network is trained as the local deterministic error model $\delta(X_{k})$ using either direct data ($\{x_{k},hc\overline{z}_{k}\}$) or indirect data (first to fourth moments of the slow variable $x_{k}$ and the autocorrelation of $x_{k}$). For the indirect data in this and the next examples, we enable the full observation of all 36 slow variables and preclude the use of all cross-terms of second to fourth moments. The direct data comprise 36000 data points, and the indirect data 154 data points. The results are presented in Figs. 4 and 5.
@@
To avoid the overfitting in Fig. 5, we calibrate a non-local error model, as discussed in Section 3.5. Compared to the results for the local error model, it can be seen in Fig. 6 that the invariant measure of the calibrated system agrees better with the true system, which indicates that the closure model $\delta(\cdot)$ in (24) with non-local effects represents a better closure for unresolved scales if there is a less clear scale separation between resolved and unresolved scales. The non-local model tends to be more flexible than the local model. Figure 5a demonstrates that the local model already achieves a good agreement in data space with the true system. It is likely that more than one possible non-local model can achieve a comparable performance of matching the indirect data. However, simulating those different non-local models for a much longer time can lead to more noticeable differences in their invariant measures, which provides a possible explanation for the larger variability within the ensemble in Fig. 6b. It should be noted that the key challenge of this example is the relatively large variability of the true closure term in Fig. 4a. Without accounting for such variability, the local and deterministic model does not correctly capture the invariant measure even in the standard regression setting with an abundance of training data as presented in Fig. 4b. In the non-local model, two local states with similar values may have noticeably different values of their neighbor states, and thus the non-local model with the local state and its neighbor states as the inputs can address some of the variability in Fig. 4a, which explains its improved performance relative to to the local models. Considering that there are three peaks in the invariant measure of the true system, the non-local model has a greater chance of having some inputs from the left or right peak than merely having inputs from the central peak. Therefore, the trained model can be more in favor of the left or right peak of the invariant measure, which provides a possible explanation for the pattern of peaks presented in Fig. 6b.
@@
We also explored learning a stochastic error model for this example. The neural networks used in the stochastic error models have 5 neurons in the first hidden layer and one neuron in the second layer. The hidden layer neurons have a sigmoid activation function, and the output layer neurons have no nonlinear activation function. For the stochastic error model with an additive noise term, 19 parameters need to be estimated. For the stochastic error model with a multiplicative noise term, 36 parameters need to be estimated. Figure 7 presents the results of the calibrated system with an additive stochastic error model. Compared to the results for deterministic error models in Figs. 5 and 6, we can see that the invariant measure of the calibrated system agrees better with the true system in Fig. 7b. We further test the stochastic error model by also learning a state-dependent diffusion coefficient. As shown in Fig. 8b, the calibrated system achieves better agreement with the invariant measure of the true system, which confirms that increased flexibility in the stochastic error model can help achieve improved predictive performance via training against indirect data. It should be noted that Figs. 6 to 8 do not display a single converged invariant measure. This arises because the calibrated parameters vary across the ensemble; as a result, we obtain a family of structural error models that all fit the data with similar accuracy. We surmise that this is caused by the fact that the indirect data contain only partial information about the invariant measure. Nonetheless, the fits are all far superior to those obtained with the local deterministic model.